# by Laurent Busé, Thomas Dedieu

### v2

@@ Abstract
Property (b) is equivalent to the vanishing of $\mathrm{Disc}(f_{s})$, and to the tangent cone to the hypersurface $\mathrm{V}(h)$ at the origin being a cone over a singular degree $s$ hypersurface (with the convention that a hypersurface of degree $s^{\prime}>s$ is a singular degree $s$ hypersurface); in other words, in the blow-up of $\mathrm{V}(h)$ at the origin, the exceptional divisor is singular.
Property (a) is implied by the vanishing of $\mathrm{Disc}(f_{d})$, as Proposition (2.2) tells us; this is equivalent to the hypersurface $\mathrm{V}(h)$ being tangent to the hyperplane at infinity $\mathrm{V}(X_{0})$.
In Definition (2.3) we discard the two factors $\mathrm{Disc}(F_{s})$ and $\mathrm{Disc}(F_{d})$; the upshot is that the vanishing of the reduced discriminant $\operatorname{redDisc}(H)$ defines a divisor which has as a dense subset the locus of those hypersurfaces $\mathrm{V}(h)$ that have a singularity off the origin $(1:0:\ldots:0)$. It turns out that this divisor is irreducible, and that the reduced discriminant is irreducible; the proof of this result will appear in [4], but we point out that this property also follows from the theory of toric discriminants (see [12, Chapter 9]) of which the reduced discriminants are specific instances.
@@
In the reduced resultant situation, the condition that the $n+1$ hypersurfaces $\mathrm{V}(h_{1}),\ldots,\mathrm{V}(h_{n+1})$ have an infinitely near additional intersection point is not divisorial, i.e., it does not correspond to an irreducible element factoring out of $\operatorname{redRes}(H_{1},\ldots,H_{n+1})$. The reason is that this condition has codimension larger than one: indeed it amounts to the $n+1$ leading terms $f_{1,s_{1}},\ldots,f_{n+1,s_{n+1}}$ having a non-trivial common zero, in other words the $n+1$ corresponding hypersurfaces in $\mathbf{P}^{n-1}$ have non-trivial intersection: this is $n+1$ hypersurfaces in $\mathbf{P}^{n-1}$, hence a condition of codimension larger than one.
In contrast, in the reduced discriminant situation the condition that the hypersurface $\mathrm{V}(h)$ has an infinitely near additional singularity, i.e., a singularity worse than an ordinary $s$-fold point, is indeed divisorial as we have seen above. It corresponds to the vanishing of $\mathrm{Disc}(F_{s})$ at the leading term $f_{s}$ of $h$.
There is an interesting connection to Milnor number. Assume that $h$ defines a hypersurface with isolated singularities, so that both its Milnor number at the origin $\mu_{0}(h)$ and its total Milnor number $\mu(h)$ are well-defined. If the only singularity of $h$ is an ordinary $s$-fold point in the origin, then $\mu(h)=\mu_{0}(f_{s})=(s-1)^{n}$. Thus, Property (b) in Paragraph (2.5) above is equivalent to the condition that $\mu_{0}(h)>(s-1)^{n}$, see [11, Theorem 1]. Therefore, it follows that if $\mathrm{Disc}(f_{d})\neq 0$, then $\mu(h)>(s-1)^{n}$ if and only $\mathrm{Disc}(f_{s})\operatorname{redDisc}(h)=0$. Pushing further, we see that if $\mathrm{Disc}(f_{d})\neq 0$ and $\mathrm{Disc}(f_{s})\neq 0$, then $\operatorname{redDisc}(h)$ vanishes if and only if $\mu(h)>\mu_{0}(h)=(s-1)^{n}$.
@@
the truncation of $F$ at order $d-s>0$ with respect to $X_{0}$. Then, setting $H=X_{1}^{s}\cdot\vphantom{H}^{\flat}\kern-1.00006ptH$ it is easy to check that $\operatorname{redDisc}_{d}^{s}(H)=\pm\mathrm{Disc}(\vphantom{H}^{\flat}\kern-1.00006ptH)$; it is therefore an irreducible polynomial of degree $2(d-s-1)$ in the coefficients of $H$ (compare with Formula (2.8.1) for the degree).
@@
(2.9.1) $(d-s)(d-s-1)$
by Proposition (2.6), as $\vphantom{H}{}^{\flat}\kern-1.00006ptH$ has degree $d-s$. Similarly, if $A_{\mathbf{Z}}$ is graded with the rule $\mathrm{weight}(U_{i})=d-i$ then the same conclusion holds (note that $d-i=(d-s)-(i-s)$).
We can generalize this following Corollary (1.11). Let $r$ be an integer and consider the grading of $A_{\mathbf{Z}}$ defined by the rule $\mathrm{weight}(U_{i})=i-s+r$ if $i\geqslant s$, and $\mathrm{weight}(U_{i})=0$ otherwise, then $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree $(d-s+2r)(d-s-1)$. In particular, if $r=s$, i.e., $\mathrm{weight}(U_{i})=i$ if $i\geqslant s$ and $\mathrm{weight}(U_{i})=0$ otherwise, we get that $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree
(2.9.2) $(d+s)(d-s-1)=d(d-1)-s(s+1).$
In fact, since $\operatorname{redDisc}_{d}^{s}(H)$ a polynomial in $U_{s},\ldots,U_{d}$ only, the weights of $U_{0},\ldots,U_{s-1}$ don’t matter. Similarly, in the grading of $A_{\mathbf{Z}}$ is defined by the rule $\mathrm{weight}(U_{i})=d-i+r$ if $i\geqslant s$ and $\mathrm{weight}(U_{i})=0$ otherwise, $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree $(d-s+2r)(d-s-1)$.
@@
Suppose that the ring $A_{\mathbf{Z}}{}$ is graded by means of the Zariski grading (1.1), i.e., $\mathrm{weight}(U_{\boldsymbol{\alpha}})=\max(\alpha_{0}-d+s,0)$. Then $\mathrm{Disc}(F)$ has valuation $s(s-1)^{n}$, and its homogeneous part $D_{0}$ in this degree satisfies the following equality in $A_{\mathbf{Z}}$:
@@
Let us point out that the three elements $\mathrm{Disc}(F_{d})$, $\mathrm{Disc}(F_{s})$, and $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$ have degree 00 with respect to the Zariski grading, while $\mathrm{Disc}(G)$ is homogeneous of degree $s(s-1)^{n}$ by Proposition (2.6.1); note that $F_{d}$ and $F_{s}$ are generic homogeneous polynomials of respective degrees $d$ and $s$ in the indeterminates $(X_{1},\ldots,X_{n})$, while $G$ is generic homogeneous of degree $s$ in $(X_{0},\ldots,X_{n})$.
@@
The only trace towards the reduced discriminant in Salmon’s work that we have found is in [19, §117], where he proves, in the above language, that for $s=2$, $\mathrm{Disc}(F)$ has valuation at least $2$.
@@
which means that it has valuation at least $2$ with respect to the Zariski grading for $s=2$.
@@
(2.11.1) $\mathrm{Res}(\partial_{1}F,\ldots,\partial_{n}F,F)$ $=\mathrm{Disc}(F)\,\mathrm{Disc}(\bar{F})$
@@
In the Zariski grading of $A_{\mathbf{Z}}{}$ as in Theorem (2.10), the discriminant $\mathrm{Disc}(F)$ has valuation $s(s-1)^{n}$, and can be written as
@@
In the grading of (2.13.1), $\mathrm{Disc}_{d}(F)$ has weight $d(d-1)^{n}$ by by Proposition (2.6). One has $G=X_{0}^{s}F_{0}+X_{0}^{s-1}F_{1}+\cdots+F_{s}$ and the coefficients of $F_{i}$ have weight $d-i=(s-i)+(d-s)$, hence $\mathrm{Disc}_{d}(G)$ has weight
@@
by Proposition (2.6) and Corollary (2.7). The coefficients of $F_{s}$ all have weight $d-s$, hence $\mathrm{Disc}_{s}(F_{s})$ has weight
@@
[…]
@@
(we introduce a rational number in the definition of $F_{2}$ to follow Salmon’s notation; this is natural in terms of the Taylor–Newton Formula of [A, Paragraph **]).
We consider the truncation of $F$ at order $d-2$ with respect to $X$, and the corresponding Zariski grading. So, $U$ has degree 2, $S$ and $T$ have degree 1, and the coefficients of the $F_{k}$’s with $k\geqslant 2$ have degree 00 (this includes the coefficients $A,B,C$). We let
@@
[…]
@@
where each “$[\ \partial_{i}F_{2}\ ]$” denotes the line containing the $n$ coefficients of $\partial_{i}F_{2}$ in the indeterminates $X_{1},\ldots,X_{n}$; from formula (2.2.5) one may easily retrace the previous computations.
Either way, the upshot is that $\mathrm{Disc}(TX_{0}X_{n}+F_{2})=-T^{2}\,\mathrm{Disc}(\bar{F}_{2})$ (note that $a(n,2)=a(n-2,2)$), and eventually we obtain the following generalized Salmon formula:
@@
[…]
We work out Salmon’s procedure in subsection 3.2, and in subsection 3.3 we show how it carries over for hypersurfaces in a projective space of arbitrary dimension. From now on, we work over an algebraically closed field $\mathbf{k}$ of characteristic 00. We use freely the theory of polarity, on which there is a recap in the previous chapter [A, Appendix A].
@@
We let $\mathbf{k}$ be an algebraically closed field of characteristic 00, and consider a smooth, degree $d$, surface $S\subseteq\mathbf{P}^{3}$ defined by a homogeneous polynomial $F\in\mathbf{k}[X,Y,Z,W]$; we set
$\mathcal{T}_{S}=\bigl{\{}(p^{\prime},p^{\prime\prime}):p^{\prime}\in S\text{ and }p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S\bigr{\}}\,\subseteq\,\mathbf{P}^{3}\times\mathbf{P}^{3}$
Let $\Phi_{\hat{p}^{\prime},\hat{p}^{\prime\prime}}(\hat{p})$ be a trihomogeneous polynomial in the sets of variables $(\hat{p}^{\prime},\hat{p}^{\prime\prime},\hat{p})\in(\mathbf{k}^{4})^{3}$, of tridegree $(\lambda,\mu,\mu)$ for some non-negative integers $\lambda,\mu$. Assume that for all $(p^{\prime},p^{\prime\prime})\in\mathcal{T}_{S}$, $p^{\prime}\neq p^{\prime\prime}$, the hypersurface $\mathrm{V}(\Phi_{\hat{p}^{\prime},\hat{p}^{\prime\prime}})\subseteq\mathbf{P}^{3}$ consists of $\mu$ planes, counted with multiplicities, all containing the line $\langle p^{\prime},p^{\prime\prime}\rangle$.
@@
[…]
@@
Now to the lemma. Projection on the first factor in $\mathbf{P}^{3}\times\mathbf{P}^{3}$ makes $\mathcal{T}_{S}$ a $\mathbf{P}^{2}$ bundle over $S$: for all $p^{\prime}\in S$, the fibre over $p^{\prime}$ is the zero locus of the linear form $\mathrm{D}_{p^{\prime}}F$ (the differential of $F$, a defining equation for $S$, at the point $p^{\prime}$; it is non-zero as we assume that $S$ is smooth). Similarly, for all $p^{\prime}\in\mathbf{P}^{3}$ the fibre $\mathrm{pr}_{1}^{-1}(p^{\prime})\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is the zero locus of the linear form $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$, which is either a hyperplane or the whole $\mathbf{P}^{3}$. We claim that for general $L^{\prime},L^{\prime\prime}$, the intersection $\mathcal{T}_{S}\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is the closure of a $\mathbf{P}^{1}$-bundle over an open subset of $S$, which implies at once that it is irreducible. Moreover, this implies that the intersection $\mathcal{T}_{S}\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is fibrewise over $S$ a transverse intersection of hyperplanes, and therefore it is reduced. Thus the claim proves the lemma. To prove the claim it suffices to show that, for general $L^{\prime},L^{\prime\prime}$, the set of points $p^{\prime}\in S$ such that the linear form $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$ is a multiple of $\mathrm{D}_{p^{\prime}}F$ has dimension 00; in particular, the intersection $\mathcal{T}_{S}\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ cannot have any irreducible component generically a $\mathbf{P}^{2}$-bundle over a curve in $S$.
@@
$\bigl{(}\partial_{X}F(p^{\prime}),\,\partial_{Y}F(p^{\prime}),\,\partial_{Z}F(p^{\prime}),\,\partial_{W}F(p^{\prime})\bigr{)}\quad\text{and}\quad\begin{pmatrix}X^{\prime\prime}\\ Y^{\prime\prime}\\ Z^{\prime\prime}\\ W^{\prime\prime}\end{pmatrix}^{\kern-2.10002pt\mathsf{T}}\cdot\begin{pmatrix}0&U_{01}&U_{02}&U_{03}\\ -U_{01}&0&U_{12}&U_{13}\\ -U_{02}&-U_{12}&0&U_{23}\\ -U_{03}&-U_{13}&-U_{23}&0\end{pmatrix}.$
Both have the line $p^{\prime}\subseteq\mathbf{k}^{4}$ in their kernel, so we consider the corresponding linear forms on the quotient $\mathbf{k}^{4}\big{/}p^{\prime}$, given by line matrices with three entries; up to a linear change of coordinates we may assume that $p^{\prime}=(0:0:0:1)$, in which case we are simply forgetting the last entry in the two above line matrices. The claim then boils down to the fact that the three $2\times 2$ minors of the $2\times 3$ matrix obtained by concatenating the two line matrices corresponding to $\mathrm{D}_{p^{\prime}}(F)$ and $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$ define a locus of codimension at least two on $S$, for general $L^{\prime},L^{\prime\prime}$. Since everything is explicit, one may observe that for a general choice of $L^{\prime},L^{\prime\prime}$, any two of these three minors have no common factor, which proves the claim. We believe it is best to let the reader write down this elementary verification on her or his own. ∎
@@
Consider $T_{p^{\prime}}=\mathrm{D}^{p}F(p^{\prime})$ as a trihomogeneous polynomial of tridegree $(d-1,0,1)$ in $(p^{\prime},p^{\prime\prime},p)$. Let $L^{\prime},L^{\prime\prime}\in\check{\mathbf{P}}^{3}$ be two generic linear forms in $p$, and consider the line $\Lambda=\mathrm{V}(L^{\prime},L^{\prime\prime})$ in $\mathbf{P}^{3}$. It intersects $\mathbf{T}_{p^{\prime}}S\cap\mathrm{V}(\Phi_{p^{\prime},p^{\prime\prime}})$ if and only if the resultant

### v3

@@
Property (b) is equivalent to the vanishing of $\mathrm{Disc}(f_{s})$, i.e., to the tangent cone of the hypersurface $\mathrm{V}(h)$ at the origin being a cone over a singular degree $s$ hypersurface (with the convention that a hypersurface of degree $s^{\prime}>s$ is a singular degree $s$ hypersurface); in other words, in the blow-up of $\mathrm{V}(h)$ at the origin, the exceptional divisor is singular. This is equivalent to the singularity of $V(h)$ at the origin being worse than an ordinary $s$-fold point, but does not imply that the blow-up of $V(h)$ is singular, see Example (2.6) below.
Property (a) is implied by the vanishing of $\mathrm{Disc}(f_{d})$, as we have seen in the proof of Proposition (2.2). The latter condition is equivalent to the hypersurface $\mathrm{V}(h)$ being tangent to the hyperplane at infinity $\mathrm{V}(X_{0})$.
In Definition (2.3) we discard the two factors $\mathrm{Disc}(F_{s})$ and $\mathrm{Disc}(F_{d})$; the upshot is that the vanishing of the reduced discriminant $\operatorname{redDisc}(H)$ defines an (irreducible444the proof of the irreducibility of the reduced discriminant will appear in [4], but this property also follows from the theory of toric discriminants (see [12, Chapter 9]), of which the reduced discriminants are particular instances. ) divisor which has as a dense subset the locus of those hypersurfaces $\mathrm{V}(h)$ that have a singularity off the origin. This implies, since smoothness is an open property, that if $\operatorname{redDisc}(H)$ vanishes at $h$, then the proper transform555here we take the following slightly abusive definition for the proper transform: let $\varepsilon:\tilde{\mathbf{P}}^{n}\to\mathbf{P}^{n}$ be the blow-up at the origin, with exceptional divisor $E$; the proper transform of $V(h)$ is the divisor $\varepsilon^{*}V(h)-sE$. If $V(h)$ has an ordinary $(s+1)$-fold at the origin, its genuine proper transform is $\hat{V}(h)=\varepsilon^{*}V(h)-(s+1)E$, which is smooth along $E$, whereas its proper transform in our sense is $\hat{V}(h)+E$, which is singular along $E$. of $V(h)$ in the blow-up of the origin is singular, in other words $V(h)$ has an additional singularity, possibly infinitely near, besides its $s$-fold point at the origin. In fact the converse holds as well, as we shall see in [4]. Thus, $\operatorname{redDisc}(H)$ vanishes at $h$ if and only $V(h)$ has an additional singularity, possibly infinitely near.
Note that we make a difference between having a singularity worse than an ordinary $s$-fold point and having an additional singularity infinitely near the ordinary $s$-fold point, see the following example and (2.8) below for more details.
(2.6) Example.
Set $n=2$ and $s=2$. In this case, $\mathrm{Disc}(F_{2})$ vanishes at $h$ if and only if $f_{2}$ is a square, equivalently, $h$ may be written as
$h=X_{0}^{d-2}\,X^{2}+X_{0}^{d-3}\,\bigl(a_{3}Y^{3}+a_{2}XY^{2}+a_{1}X^{2}Y+a_{0}X^{3}\bigr)+\,\cdots$
after a suitable change of coordinates $(X,Y)=\phi(X_{1},X_{2})$. Let $\varepsilon:\tilde{\mathbf{P}}^{2}\to\mathbf{P}^{2}$ be the blow-up at the origin. The relevant place to study the proper transform of $V(h)$ over the origin $(1\!:\!0\!:\!0)\in\mathbf{P}^{2}$ is the affine plane with coordinates $(U,Y)$ in $\tilde{\mathbf{P}}^{2}$ which dominates the affine plane $(X_{0}=1)$ in $\mathbf{P}^{2}$, and such that $\varepsilon^{*}(X,Y)=(UY,Y)$. In this affine chart, the pull-back $\varepsilon^{*}h$ is given by
$U^{2}Y^{2}+\bigl(a_{3}Y^{3}+a_{2}UY^{3}+a_{1}U^{2}Y^{3}+a_{0}U^{3}Y^{3}\bigr)+\,\cdots$
and the proper transform of $V(h)$ is defined by the equation obtained by factoring out $Y^{2}$ from $\varepsilon^{*}h$, i.e.,
$U^{2}+\bigl(a_{3}Y+a_{2}UY+a_{1}U^{2}Y+a_{0}U^{3}Y\bigr)+\,\cdots$
This shows that the proper transform of $V(h)$ is singular at some point lying on the exceptional divisor of $\varepsilon$ (defined by $Y=0$ in this affine chart) if and only if $a_{3}=0$.
The generic $h$ such that $\mathrm{Disc}(F_{2})$ vanishes defines a plane curve with an ordinary cusp at the origin (local equation $x^{2}+y^{3}=0)$; the proper transform of this curve in the blow-up at the origin is smooth (and tangent to the exceptional divisor), hence $\operatorname{redDisc}(H)$ does not vanish at $h$ even though $V(h)$ has a singularity worse than an ordinary double point at the origin. On the other hand, the generic $h$ such that $V(h)$ has an additional singularity infinitely near the origin (which implies that $\operatorname{redDisc}(H)$ vanishes) defines a plane curve with a tacnode at the origin (local equation $x^{2}+y^{4}=0$, two smooth local branches tangent to each other); the proper transform of such a curve has an ordinary double point located on the exceptional divisor.
@@
In the reduced resultant situation, the condition that the $n+1$ hypersurfaces $\mathrm{V}(h_{1}),\ldots,\mathrm{V}(h_{n+1})$ have an infinitely near additional intersection point is not divisorial, i.e., it does not correspond to an irreducible element factoring out of $\operatorname{redRes}(H_{1},\ldots,H_{n+1})$. The reason is that this condition has codimension larger than one: indeed it amounts to the $n+1$ leading terms $f_{1,s_{1}},\ldots,f_{n+1,s_{n+1}}$ having a non-trivial common zero, in other words the $n+1$ corresponding hypersurfaces in $\mathbf{P}^{n-1}$ have non-trivial intersection: these are $n+1$ hypersurfaces in $\mathbf{P}^{n-1}$, hence the condition has codimension larger than one.
In contrast, in the reduced discriminant situation the condition that the hypersurface $\mathrm{V}(h)$ has a singularity worse than an ordinary $s$-fold point is divisorial, as it corresponds to the vanishing of $\mathrm{Disc}(F_{s})$ at the leading term $f_{s}$ of $h$. The vanishing of $\operatorname{redDisc}(H)$ at $h$ is not implied by the latter condition, but by the stronger condition that $V(h)$ has an additional singularity infinitely near its $s$-fold point.
We may use the Milnor number (see [I, Section **] for the definition) to give a precise meaning to the condition of having a singlarity worse than an ordinary $s$-fold point. Assume that $h$ defines a hypersurface with an isolated singularity at the origin, so that its Milnor number at the origin $\mu_{0}(h)$ is well-defined. If the singularity of $h$ at the origin is an ordinary $s$-fold point, then $\mu_{0}(h)=(s-1)^{n}$, and condition (b) in Paragraph (2.5) above is equivalent to the condition that $\mu_{0}(h)>(s-1)^{n}$ by [11, Theorem 1].
@@
the truncation of $F$ at order ds>0d-s>0 with respect to $X_{0}$. Then, setting $H=X_{1}^{s}\cdot\vphantom{H}^{\flat}\kern-1.00006ptH$ it is easy to check that $\operatorname{redDisc}_{d}^{s}(H)=\pm\mathrm{Disc}(\vphantom{H}^{\flat}\kern-1.00006ptH)$; it is therefore an irreducible polynomial of degree 2(ds1)2(d-s-1) in the coefficients of $H$ (compare with Formula (2.9.1) for the degree).
@@
(2.10.1) (ds)(ds1)(d-s)(d-s-1)
by Proposition (2.6), as $\vphantom{H}{}^{\flat}\kern-1.00006ptH$ has degree dsd-s. Similarly, if $A_{\mathbf{Z}}$ is graded with the rule $\mathrm{weight}(U_{i})=d-i$ then the same conclusion holds (note that di=(ds)(is)d-i=(d-s)-(i-s)).
We can generalize this following Corollary (1.12). Let $r$ be an integer and consider the grading of $A_{\mathbf{Z}}$ defined by the rule $\mathrm{weight}(U_{i})=i-s+r$ if $i\geqslant s$, and $\mathrm{weight}(U_{i})=0$ otherwise, then $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree (ds+2r)(ds1)(d-s+2r)(d-s-1). In particular, if $r=s$, i.e., $\mathrm{weight}(U_{i})=i$ if $i\geqslant s$ and $\mathrm{weight}(U_{i})=0$ otherwise, we get that $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree
(2.10.2) (d+s)(ds1)=d(d1)s(s+1).(d+s)(d-s-1)=d(d-1)-s(s+1).
In fact, since $\operatorname{redDisc}_{d}^{s}(H)$ a polynomial in $U_{s},\ldots,U_{d}$ only, the weights of $U_{0},\ldots,U_{s-1}$ don’t matter. Similarly, in the grading of $A_{\mathbf{Z}}$ is defined by the rule $\mathrm{weight}(U_{i})=d-i+r$ if $i\geqslant s$ and $\mathrm{weight}(U_{i})=0$ otherwise, $\operatorname{redDisc}_{d}^{s}(H)$ is homogeneous of degree (ds+2r)(ds1)(d-s+2r)(d-s-1).
@@
Suppose that the ring $A_{\mathbf{Z}}{}$ is graded by means of the Zariski grading (1.1), i.e., $\mathrm{weight}(U_{\boldsymbol{\alpha}})=\max(\alpha_{0}-d+s,0)$. Then $\mathrm{Disc}(F)$ has order $s(s-1)^{n}$, and its homogeneous part $D_{0}$ in this degree satisfies the following equality in $A_{\mathbf{Z}}$:
@@
Let us point out that the three elements $\mathrm{Disc}(F_{d})$, $\mathrm{Disc}(F_{s})$, and $\operatorname{redRes}(\partial_{1}H,\ldots,\partial_{n}H,H)$ have degree 0 with respect to the Zariski grading, while $\mathrm{Disc}(G)$ is homogeneous of degree $s(s-1)^{n}$ by (2.6.1); note that $F_{d}$ and $F_{s}$ are generic homogeneous polynomials of respective degrees $d$ and $s$ in the indeterminates $(X_{1},\ldots,X_{n})$, while $G$ is generic homogeneous of degree $s$ in $(X_{0},\ldots,X_{n})$.
@@
The only trace towards the reduced discriminant in Salmon’s work that we have found is in [19, §117], where he proves, in the above language, that for $s=2$, $\mathrm{Disc}(F)$ has order at least $2$.
@@
which means that it has order at least $2$ with respect to the Zariski grading for $s=2$.
@@
($2.1$) $\mathrm{Res}(\partial_{1}F,\ldots,\partial_{n}F,F)$ $=\mathrm{Disc}(F)\,\mathrm{Disc}(\bar{F})$
@@
In the Zariski grading of $A_{\mathbf{Z}}{}$ as in Theorem (2.11), the discriminant $\mathrm{Disc}(F)$ has order $s(s-1)^{n}$, and can be written as
@@
In the grading of (2.14.1), $\mathrm{Disc}_{d}(F)$ has weight $d(d-1)^{n}$ by Proposition (2.6). One has $G=X_{0}^{s}F_{0}+X_{0}^{s-1}F_{1}+\cdots+F_{s}$ and the coefficients of $F_{i}$ have weight di=(si)+(ds)d-i=(s-i)+(d-s), hence $\mathrm{Disc}_{d}(G)$ has weight
@@
by Proposition (2.6) and Corollary (2.7). The coefficients of $F_{s}$ all have weight dsd-s, hence $\mathrm{Disc}_{s}(F_{s})$ has weight
@@
We shall now see that Salmon’s formula (a) is a particular case of the decomposition formula given in Corollary (2.13). We will then be able to generalize it to the case of a hypersurface in arbitrary dimension.
@@
(we introduce a rational number in the definition of $F_{2}$ to follow Salmon’s notation; this is natural in terms of the Taylor–Newton Formula of [II, Paragraph **]).
We consider the truncation of $F$ at order d2d-2 with respect to $X$, and the corresponding Zariski grading. So, $U$ has degree 2, $S$ and $T$ have degree 1, and the coefficients of the $F_{k}$’s with $k\geqslant 2$ have degree 0 (this includes the coefficients $A,B,C$). We let
@@
this is Salmon’s Formula (a), with the normalization factor $\frac{1}{2}$, and with $\Phi=\operatorname{redDisc}(H)$.
@@
[…]
Either way, the upshot is that $\mathrm{Disc}(TX_{0}X_{n}+F_{2})=-T^{2}\,\mathrm{Disc}(\bar{F}_{2})$ (note that a(n,2)=a(n2,2)a(n,2)=a(n-2,2)), and eventually we obtain the following generalized Salmon formula:
@@
Indeed, the line $\langle p^{\prime},p^{\prime\prime}\rangle^{\perp}$ is contained in the plane $(p^{\prime})^{\perp}$, hence tangent to $S^{\vee}$ at the point $(\mathbf{T}_{p^{\prime}}S)^{\perp}\in S^{\vee}$. If the plane $\mathbf{T}_{p^{\prime}}S$ is plainly bitangent to $S$, then $(\mathbf{T}_{p^{\prime}}S)^{\perp}$ is a general point on the ordinary double curve of $S^{\vee}$, and the line $\langle p^{\prime},p^{\prime\prime}\rangle^{\perp}$ is tangent to one of the two transverse sheets of $S^{\vee}$ at $(\mathbf{T}_{p^{\prime}}S)^{\perp}$. The idea is then to determine first the conditions on $p^{\prime}$ for $\mathbf{T}_{p^{\prime}}S$ to count with multiplicity greater than $2$ in $\langle p^{\prime},p^{\prime\prime}\rangle^{\perp}$, and then to sort out the various corresponding geometric situations. A key element to carry this out is the famous formula (a); another one is the elimination procedure (3.2).
We work out Salmon’s procedure in subsection 3.2, and in subsection 3.3 we show how it carries over for hypersurfaces in a projective space of arbitrary dimension. From now on, we work over an algebraically closed field $\mathbf{k}$ of characteristic 0. We use freely the theory of polarity, on which there is a recap in the previous chapter [II, Appendix A].
@@
Let $\mathbf{k}$ be an algebraically closed field of characteristic 0, and consider a smooth, degree $d$, surface $S\subseteq\mathbf{P}^{3}$ defined by a homogeneous polynomial $F\in\mathbf{k}[X,Y,Z,W]$; set
$\mathcal{T}_{S}=\bigl\{(p^{\prime},p^{\prime\prime}):p^{\prime}\in S\text{ and }p^{\prime\prime}\in\mathbf{T}_{p^{\prime}}S\bigr\}\,\subseteq\,\mathbf{P}^{3}\times\mathbf{P}^{3}.$
[…]
@@
$D_{L^{\prime},L^{\prime\prime}}(p^{\prime},p^{\prime\prime})\overset{\text{def.}}{=}\begin{vmatrix}L^{\prime}(\hat{p}^{\prime})&L^{\prime\prime}(\hat{p}^{\prime})\\ L^{\prime}(\hat{p}^{\prime\prime})&L^{\prime\prime}(\hat{p}^{\prime\prime})\end{vmatrix}$
@@
Now to the lemma. Projection on the first factor in $\mathbf{P}^{3}\times\mathbf{P}^{3}$ makes $\mathcal{T}_{S}$ a $\mathbf{P}^{2}$ bundle over $S$: for all $p^{\prime}\in S$, the fibre over $p^{\prime}$ is the zero locus of the linear form $\mathrm{D}_{p^{\prime}}F$ (the differential of $F$, a defining equation for $S$, at the point $p^{\prime}$; it is non-zero as we assume that $S$ is smooth). Similarly, for all $p^{\prime}\in\mathbf{P}^{3}$ the fibre $\mathrm{pr}_{1}^{-1}(p^{\prime})\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is the zero locus of the linear form $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$, which is either a hyperplane or the whole $\mathbf{P}^{3}$. We claim that for general $L^{\prime},L^{\prime\prime}$, the intersection $\mathcal{T}_{S}\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is the closure of a $\mathbf{P}^{1}$-bundle over an open subset of $S$, which implies at once that it is irreducible. Moreover, this implies that the intersection $\mathcal{T}_{S}\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ is fibrewise over $S$ a transverse intersection of hyperplanes, and therefore it is reduced. Thus the claim proves the lemma. To prove the claim it suffices to show that, for general $L^{\prime},L^{\prime\prime}$, the set of points $p^{\prime}\in S$ such that the linear form $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$ is a multiple of $\mathrm{D}_{p^{\prime}}F$ has dimension 0; in particular, the intersection $\mathcal{T}_{S}\cap\mathrm{V}(D_{L^{\prime},L^{\prime\prime}})$ cannot have any irreducible component generically a $\mathbf{P}^{2}$-bundle over a curve in $S$.
@@
$\bigl(\partial_{X}F(p^{\prime}),\,\partial_{Y}F(p^{\prime}),\,\partial_{Z}F(p^{\prime}),\,\partial_{W}F(p^{\prime})\bigr)\quad\text{and}\quad\begin{pmatrix}X^{\prime\prime}\\ Y^{\prime\prime}\\ Z^{\prime\prime}\\ W^{\prime\prime}\end{pmatrix}^{\kern-2.45836pt\mathsf{T}}\cdot\begin{pmatrix}0&U_{01}&U_{02}&U_{03}\\ -U_{01}&0&U_{12}&U_{13}\\ -U_{02}&-U_{12}&0&U_{23}\\ -U_{03}&-U_{13}&-U_{23}&0\end{pmatrix}$.
Both have the line $p^{\prime}\subseteq\mathbf{k}^{4}$ in their kernel, so we consider the corresponding linear forms on the quotient $\mathbf{k}^{4}\big/p^{\prime}$, given by line matrices with three entries; up to a linear change of coordinates we may assume that $p^{\prime}=(0:0:0:1)$, in which case we are simply forgetting the last entry in the two above line matrices. The claim then boils down to the fact that the three $2\times 2$ minors of the $2\times 3$ matrix obtained by concatenating the two line matrices corresponding to $\mathrm{D}_{p^{\prime}}(F)$ and $D_{L^{\prime},L^{\prime\prime}}(p^{\prime},\,\_\,)$ define a locus of codimension at least two on $S$, for general $L^{\prime},L^{\prime\prime}$. Since everything is explicit, one may observe that for a general choice of $L^{\prime},L^{\prime\prime}$, any two of these three minors have no common factor, which proves the claim. ∎
@@
Consider $T_{p^{\prime}}=\mathrm{D}^{p}F(p^{\prime})$ as a trihomogeneous polynomial of tridegree (d1,0,1)(d-1,0,1) in $(p^{\prime},p^{\prime\prime},p)$. Let $L^{\prime},L^{\prime\prime}\in\check{\mathbf{P}}^{3}$ be two generic linear forms in $p$, and consider the line $\Lambda=\mathrm{V}(L^{\prime},L^{\prime\prime})$ in $\mathbf{P}^{3}$. It intersects $\mathbf{T}_{p^{\prime}}S\cap\mathrm{V}(\Phi_{p^{\prime},p^{\prime\prime}})$ if and only if the resultant